Table 13-7
The Flying Elvis Copter Rides

Quantity Total
Cost
Fixed
Cost
Variable
Cost
Marginal
Cost
Average
Fixed
Cost
Average
Variable
Cost
Average
Total
Cost
0 $50 $50 $0 -- -- -- --
1 $150 A B C D E F
2 G H I $120 J K L
3 M N O P Q $120 R



Refer to Table 13-7. What is the value of G?

Group of answer choices

$120

$220

$270

$30
Answer #1

Answer
the total cost of n units =total cost of n-1 units +marginal cost of n th unit
TC(n)=TC(n-1)+MC(n)
TC(2)=150+120=$270

the total cost is called G which is $270
Option 3
Answer #2
Q. What is the value of E?
Output TC FC VC MC AFC AVC ATC
0 50 50 - - - - -
1 150 50 100 100 50 100 150
2 270 50 220 120 25 110 135
3 410 50 360 140 16.66 120 136.66

Value of E is $100 Option A

Formula used

TC = TVC+ TFC

AVC = TVC/Q

AFC= FC/Q

AC= TC/Q

MC= TC(q2) - TC(q1)